# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes

### v1

@@ 3.1 Overall performance: downstream utility
We assess overall performance based on Utility of the augmented data, which we evaluate in terms of AUC on the real $D_{\mathrm{test}}$, when using four different types of downstream models (see Appendix B). This setup mirrors the widely adopted Train-on-synthetic-Test-on-real (TSTR) (Esteban et al., 2017). Additionally, we compare the performance to training on the small $D_{\mathrm{train}}$, as well as training on the large held-out $D_{\mathrm{oracle}}$, the latter serving as an upper bound.
GPT-4 + Curation has best overall performance. Table 3 shows the performance of the proposed CLLM  (GPT- 4 and GPT-3.5) against baselines — both with and without our curation mechanism. We find that the GPT-4 + Curation variant of CLLM  outperforms baselines in almost all settings ($20/28$). Interestingly, its performance is close to or even exceeds the performance of $D_{\mathrm{oracle}}$. Table 4 further shows that GPT-4 + Curation ranks first on average among all the generative methods.
Sample size sensitivity. We now investigate the performance gains of CLLM  as we vary the number of samples $n$ in $D_{\mathrm{train}}$, in Table 3 and Table 4. Performance improvements and high ranking across datasets for CLLM (GPT-4+Curation) are especially noticeable in the ultra low-data regime (i.e. $n<100$). In this regime, the limited size of $D_{\mathrm{train}}$ severely constrains the other baseline methods. In contrast, as illustrated in Sec. 2.1, CLLM  can leverage GPT-4’s prior knowledge to extrapolate beyond the small $D_{\mathrm{train}}$, thereby improving downstream performance. As expected, the performance gap between CLLM  and other methods decreases as the size of $D_{\mathrm{train}}$ grows (e.g. $n=200$), where sufficient training data helps other generators achieve good performance. We further decouple the prior knowledge of the LLM and the number of in-context samples in Appendix C.3, showing the importance of the in-context samples to guide the LLM’s generation.
@@
Curation generally helps all generative models. Our curation mechanism consistently benefits all generative models for the different $n$. It ensures only high quality samples are retained, which is crucial for good data augmentation and downstream performance and has been overlooked in previous works. This explains why the combination of the best generative model and curation, which is CLLM, gives the best results and highest rankings in the low-data regime (e.g. $n=20$).
@@
Remark on ICL versus fine-tuning. Our results in Table 3 and Table 4 indicate that ICL is better than fine-tuning (GReaT baseline) in the low-data regime. This highlights the difficulty of fine-tuning in this regime, where it is easy to overfit to $D_{\mathrm{train}}$. As we increase the number of samples, this baseline coupled with curation improves toward the level of CLLM  (GPT-4).
@@ 3.2 Hardness: a proxy signal to flag poor quality synthetic datasets
Having a systematic way to assess datasets generated by LLMs like GPT-4 is important because their black-box nature provides little control on their generation quality. This contrasts conventional generators for which training loss is an exploitable signal. Hence, we ask: could we have a signal to identify a potential problematic dataset generated by GPT-4 without an exhaustive manual review? For example, GPT-4 produced low-quality synthetic data for the Adult dataset (across the different sample sizes) resulting in poor downstream performance. While curation improves it, downstream performance is still suboptimal. Addressing this question is important, since datasets are rarely created by the ML model builder in real-world ML workflows, but rather by specialist data teams or data owners (Gebru et al., 2021; Sambasivan et al., 2021; Goncalves et al., 2020). Hence, having a signal to preemptively flag a potentially suboptimal generated dataset spares investment in both storing the subpar data and/or training a model likely to underperform on real data.
$D_{\mathrm{syn}}$ should intuitively be considered imperfect if curation discards many of its samples, since the number of discarded samples measures the quality of samples with respect to the small but gold-standard $D_{\mathrm{train}}$. Hence, we investigate the relationship between test performance (AUC) and the proportion of samples discarded by the curation. Fig. 6, where each point is a synthetic dataset generated by GPT-4 (e.g. Adult, Compas), shows a strong negative linear relationship between these two quantities. This holds across the different $n$ with slopes fairly stable around $-1.4$. This relationship corroborates the poor quality of the dataset generated by GPT-4 on the Adult dataset, providing a useful proxy that $D_{\mathrm{syn}}$ is unlikely to lead to good downstream performance.
@@ 4 Discussion
We introduce CLLM, an approach for data augmentation in the ultra low-data setting. CLLM  exploits the prior knowledge of LLMs along with curation for improved downstream performance. As empirically shown, CLLM  outperforms traditional generative models—most noticeably on underrepresented subgroups, for which data augmentation is of utmost importance. CLLM  is grounded in the ICL capability of LLMs, and benefits from its simplicity. We studied GPT-3.5 and GPT-4 as backbones for CLLM. The cost of the API access pose limitations, e.g. on wide accessibility, on knowing which data was used for training the models, and on understanding the LLM’s output better. Using smaller and open LLMs could overcome these limitations, though this could come with a reduction in performance. We leave this as a promising direction for future work. Further improvements may be achieved through different tuning and prompting of the LLM, as shown in different domains (Meng et al., 2023; Liu et al., 2023). Improving LLM tuning and prompting is beyond the scope of our work, but we regard this as a promising avenue for future work.
Ethics. In this work, we evaluate CLLM  using multiple real-world datasets. The private datasets are de-identified and used in accordance with the guidance of the respective data providers. We follow recommendations to use the Azure OpenAI service when using GPT-4 and GPT-3.5 models, where via the agreement we ensure the medical data is not sent for human review or stored, hence respecting the guidelines given by the dataset providers. LLMs may make errors and may reflect or exacerbate societal biases that are present in their data (Li et al., 2023). Though the curation in CLLM improves synthetic data quality, it does not directly aim to remove biases. The quality and fairness of generated data should always be evaluated. More research into LLM bias is required before methods like CLLM should be applied to real-world sensitive settings like healthcare and finance.
@@
Finally, increasing access to ML across regions, domains and societies is also about more than just technology. We believe broader engagement and discussion with various stakeholders is crucial to responsibly expand ML access, thereby realizing the benefits of ML in an equitable way.
@@
Data scarcity and computational limitations are deterrents for developing ML. These challenges should inspire cutting-edge ML research (De-Arteaga et al., 2018). We believe CLLM  takes a step in this direction toward improving the use of ML in low-data settings, across society (e.g. underrepresented subgroups (Suresh and Guttag, 2021)), domains (e.g. healthcare (Alami et al., 2020; Owoyemi et al., 2020)) and regions (e.g. LMICs).
@@
Appendix: Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes
@@ Appendix A Extended related work
Transfer learning [Pan and Yang, 2009], domain adaptation [Farahani et al., 2021], and few-shot learning [Wang et al., 2020] employ additional data resources or rely on specific task-related assumptions to improve learning performance. These methods exploit large labeled data from a source domain, unlabeled data in a target domain, or leverage knowledge from related tasks respectively. For example, Levin et al. [2022] and Jin and Ucar [2023] use models trained on labeled data from a source domain, while Ruiz et al. [2023] and Margeloiu et al. [2022] leverage knowledge-graphs. This is in contrast to our setting, considered of learning with limited data, which must function with whatever scarce labeled data it has, without making any assumptions about the availability of additional data or tasks.
Active learning [Settles, 2009] and semi-supervised learning [van Engelen and Hoos, 2019, Chapelle et al., 2006] also operate under the premise of having access to plentiful unlabeled data and the capacity to interactively query labels. However, in our setting, considered learning with limited data does not inherently assume such capabilities, focusing instead on limited labeled data only.
@@
Furthermore, active learning primarily focuses on the iterative process of selecting data samples that, when labeled, are expected to most significantly improve the model’s performance. This selection is typically based on criteria such as uncertainty sampling which focuses on epistemic uncertainty [Mussmann and Liang, 2018, Houlsby et al., 2011, Kirsch et al., 2019, Nguyen et al., 2022]. The primary objective is to minimize labeling effort while maximizing the model’s learning efficiency. Additionally, active learning would aims to label instances based on epistemic uncertainty where the model struggles to make accurate predictions, yet the samples themselves are correct. In contrast, CLLM leverage training dynamics based on aleatoric uncertainty and confidence and is designed to discard samples that might jeopardize the downstream accuracy. These samples can be considered to have inherent issues or are erroneous, such as being ”mislabeled”. To summarize, in active learning, epistemic uncertainty is used to identify data points that, if labeled, would yield the most significant insights for model training. In our approach, they serve to identify and exclude/filter data points that could potentially deteriorate the model’s performance.
@@
Self-supervised learning [Liu et al., 2021] leverages large amounts of unlabeled data to learn useful representations for downstream tasks. However, in our setting, considered learning with limited data does not inherently assume such access to vast amounts of unlabeled data.
@@
Data-centric AI. Ensuring high data quality is a critical but often overlooked problem in ML, where the focus is optimizing models [Sambasivan et al., 2021]. Even when it is considered, the process of assessing datasets is adhoc or artisanal [Seedat et al., 2022b]. However, the recent push of data-centric AI [Liang et al., 2022, Polyzotis and Zaharia, 2021, Zha et al., 2023] aims to develop systematic tools to curate existing datasets. Our work contributes to this nascent body of work [Seedat et al., 2023] – presenting CLLM, which, to the best of our knowledge, is the first systematic data-centric framework looking at how we can tailor synthetic datasets) rather than real datasets) to downstream task use with data curation.
@@
Why Data Augmentation? Data augmentation is a flexible approach to address the ultra low-data regime. An alternative might be to resort to a pretrained black-box model for classification, which could be for example via in-context learning for classification [Dong et al., 2022]. However, such a solution is inadequate for several reasons, many of which would prevent real-world utility (e.g. in LMICs):
@@ B.1 Datasets
Adult Income [Asuncion and Newman, 2007] 30k 12 Finance
Compas [Angwin et al., 2016] 5k 13 Criminal justice
Covid-19 [Baqui et al., 2020] 7k 29 Healthcare/Medicine
CUTRACT Prostate [PCUK, 2019] 2k 12 Healthcare/Medicine
Drug [Fehrman et al., 2017] 2k 27 Healthcare/Medicine
MAGGIC [Pocock et al., 2013] 41k 29 Healthcare/Medicine
SEER Prostate [Duggan et al., 2016] 20k 12 Healthcare/Medicine
@@
    Open-source: Adult, Drug and Compas are widely used open-source datasets used in the tabular data literature. Adult and Drug are both UCI datasets that have been used in many papers, while Compas is part of OpenML Vanschoren et al. [2013]. Our reason for selecting them is that, despite them being open-source, they are highly reflective of domains in which we might be unable to collect many samples — hence in reality would often be in an ultra-low data regime.
@@
GPT-4 and GPT-3.5 We access GPT-4 [OpenAI, 2023] and GPT-3.5-Turbo [Brown et al., 2020] through the API. We use a temperature of $0.9$.
GReaT. GReaT Borisov et al. [2023] is a generative model which fine-tunes an LLM based on a training set. We use the implementation provided by authors.
Generative model based approaches. For the other baselines used in 3.1, we use the library SynthCity [Qian et al., 2023], using the defaults. We detail each next.
  • TVAE: this is a conditional Variational Auto Encoder (VAE) for tabular data and is based on Xu et al. [2019]
  • CTGAN: A conditional generative adversarial network which can handle tabular data and is based on Xu et al. [2019]
@@
  • TabDDPM: A diffusion model that can be universally applied to any tabular dataset, handles any type of feature and is based on Kotelnikov et al. [2022]
Traditional Data Augmentation. We use SMOTE [Chawla et al., 2002] which augments data by considering nearest neighbors and performing linear interpolations. We use the implementation provided by Lemaître et al. [2017], and set the number of neighbors $k$ to $5$.

### v3

@@ 3.1 Overall performance: downstream utility
We assess overall performance based on Utility of the augmented data, which we evaluate in terms of AUC on the real $D_{\mathrm{test}}$, when using four different types of downstream models (see Appendix B). This setup mirrors the widely adopted Train-on-synthetic-Test-on-real (TSTR) (Esteban et al., 2017). Additionally, we compare the performance to training on the small $D_{\mathrm{train}}$, as well as training on the large held-out $D_{\mathrm{oracle}}$, the latter serving as an upper bound. GPT-4 + Curation has best overall performance. Table 3 shows the performance of the proposed CLLM  (GPT- 4 and GPT-3.5) vs baselines — both with and without our curation mechanism. We find that the GPT-4 + Curation variant of CLLM  outperforms baselines in almost all settings ($20/28$). Interestingly, its performance is close to or even exceeds the performance of $D_{\mathrm{oracle}}$. Table 4 further shows that GPT-4 + Curation ranks first on average vs all generative methods.
Sample size sensitivity. We now investigate the performance gains of CLLM  as we vary the number of samples $n$ in $D_{\mathrm{train}}$, in Table 3 and Table 4. Performance improvements and high ranking across datasets for CLLM (GPT-4+Curation) are especially noticeable in the low-data regime (i.e. $n<100$). In this regime, the limited size of $D_{\mathrm{train}}$ severely constrains the other baseline methods. In contrast, as illustrated in Sec. 2.1, CLLM  can leverage GPT-4’s prior knowledge to extrapolate beyond the small $D_{\mathrm{train}}$, thereby improving downstream performance. As expected, the performance gap between CLLM  and other methods decreases as the size of $D_{\mathrm{train}}$ grows (e.g. $n=200$), where sufficient training data helps other generators achieve good performance. We further decouple the prior knowledge of the LLM and the number of in-context samples in Appendix C.1, showing the importance of the in-context samples to guide the LLM’s generation.
@@
Curation generally helps all generative models. Our curation mechanism consistently benefits all generative models for the different $n$. It ensures that only high-quality samples are retained, which is crucial for good data augmentation and downstream performance and has been overlooked in previous works. This explains why the combination of the best generative model and curation (CLLM) gives the best results and highest rankings in the low-data regime (e.g. $n=20$). In addition to GPT-4 and GPT-3.5, we show the versatility of our proposed curation mechanism to provide benefit with other open-source LLM backbones, including Mistral-7b (Jiang et al., 2023), LLAMA-2 (Touvron et al., 2023) and Mixtral (Jiang et al., 2024) (cf. Appendix C.2).
@@
Remark on ICL versus fine-tuning. Our results in Table 3 and Table 4 indicate that ICL is better than fine-tuning (GReaT baseline) in the low-data regime. This highlights the difficulty of fine-tuning in this regime, where it is easy to overfit to $D_{\mathrm{train}}$. As we increase the number of samples, this baseline, coupled with curation, improves to the level of CLLM  (GPT-4).
@@ 3.2 Hardness: a proxy signal to flag poor quality synthetic datasets
Having a systematic way to assess datasets generated by LLMs like GPT-4 is important because their black-box nature provides little control on their generation quality. This contrasts conventional generators for which training loss is an exploitable signal. Hence, we ask: could we have a signal to identify a potential problematic dataset generated by an LLM without an exhaustive manual review? For example, GPT-4 produced low-quality synthetic data for the Adult dataset (across the different sample sizes) resulting in poor downstream performance. While curation improves it, downstream performance is still suboptimal.
@@
Addressing this question is important, since datasets are rarely created by the ML model builder in real-world ML workflows, but rather by specialist data teams or data owners (Gebru et al., 2021; Sambasivan et al., 2021; Goncalves et al., 2020). Thus, having a signal to preemptively flag a potentially suboptimal generated dataset spares investment in both storing the subpar data and/or training a model likely to underperform on real data.
To address this, we posit that $D_{\mathrm{syn}}$ should intuitively be considered imperfect if curation discards many of its samples, since the number of discarded samples measures the quality of samples with respect to the small but gold-standard $D_{\mathrm{train}}$.
Hence, we investigate the relationship between test performance (AUC) and the proportion of samples discarded by curation. Fig. 6, where each point is a synthetic dataset generated by GPT-4 (e.g. Adult, Compas), shows a strong negative linear relationship between these two quantities. This holds across the different $n$ with slopes fairly stable around $-1.4$. This relationship corroborates the poor quality of the dataset generated by GPT-4 on the Adult dataset, providing a useful proxy that $D_{\mathrm{syn}}$ is unlikely to lead to good downstream performance.
@@ 4 Discussion
We introduce CLLM, an approach for data augmentation in the low-data setting. CLLM  exploits the prior knowledge of LLMs along with our curation mechanism for improved downstream performance.
@@
As empirically shown, CLLM  outperforms traditional generative models—most noticeably on underrepresented subgroups, for which data augmentation is of utmost importance. CLLM  is grounded in the ICL capability of LLMs. Further improvements may be achieved through different tuning and prompting of the LLM, as shown in different domains (Meng et al., 2023; Liu et al., 2023). Improving LLM tuning and prompting is beyond the scope of our work, but we regard this as a promising avenue for future work.
While the key contribution of this work is the curation of LLM outputs, overall downstream performance gains are still fundamentally tied to the LLM backbone. Specifically, a practical consideration is that using less parameterized LLMs leads to poorer uncurated data. That said, our curation mechanism naturally addresses this aspect and improves downstream performance (see Appendix C.2).
@@
Finally, while CLLM  addresses the data logjam issue, increasing access to ML across regions, domains and societies is also about more than just technology. We believe broader engagement and discussion with various stakeholders is crucial to responsibly expand ML access, thereby realizing the benefits of ML in an equitable way.
Impact statement
@@
Data scarcity and computational limitations are deterrents for developing ML. These challenges should inspire cutting-edge ML research (De-Arteaga et al., 2018). We believe CLLM  takes a step in this direction toward improving the use of ML in low-data settings, across society (e.g. underrepresented subgroups (Suresh & Guttag, 2021)), domains (e.g. healthcare (Alami et al., 2020; Owoyemi et al., 2020)) and regions (e.g. LMICs).
However, with that in mind, LLMs may make errors and may reflect or exacerbate societal biases that are present in their data (Li et al., 2023). Though the curation in CLLM improves synthetic data quality, it does not directly aim to remove biases. The quality and fairness of generated data should always be evaluated. We believe broader engagement and discussion with various stakeholders is required before methods like CLLM should be applied to real-world sensitive settings like healthcare and finance, as well as more research into LLM bias and potential mitigation strategies. We provide a more detailed discussion in Section C.12.
Additionally, in this work, we evaluate CLLM  using multiple real-world datasets. The private datasets are de-identified and used in accordance with the guidance of the respective data providers. We follow recommendations to use the Azure OpenAI service when using GPT-4 and GPT-3.5 models, where via the agreement we ensure the medical data is not sent for human review or stored, hence respecting the guidelines given by the dataset providers.
Acknowledgments
The authors are grateful to Fergus Imrie, Andrew Rashbass and the anonymous ICML reviewers for their useful comments and feedback. Nabeel Seedat is supported by the Cystic Fibrosis Trust, Nicolas Huynh by Illumina, and Boris van Breugel by the Office of Naval Research UK. This work was supported by Microsoft’s Accelerate Foundation Models Academic Research initiative.
@@ Acknowledgments
Appendix: Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in low-data regimes
@@ Appendix A Extended Related Work
Transfer learning (Pan & Yang, 2009), domain adaptation (Farahani et al., 2021), and few-shot learning (Wang et al., 2020) employ additional data resources or rely on specific task-related assumptions to improve learning performance. These methods exploit large labeled data from a source domain, unlabeled data in a target domain, or leverage knowledge from related tasks respectively. For example, (Levin et al., 2022) and (Jin & Ucar, 2023) use models trained on labeled data from a source domain, while (Ruiz et al., 2023) and (Margeloiu et al., 2022) leverage knowledge-graphs. This is in contrast to our setting, considered of learning with limited data, which must function with whatever scarce labeled data it has, without making any assumptions about the availability of additional data or tasks.
Detailed contrast between CLLM  and transfer learning / meta-learning / few-shot learning.
In addition to the above, we emphasize below three dimensions along which CLLM  differs from the transfer learning and meta-learning literature (which permit to do few-shot learning). Specifically, we highlight three specific dimensions which explain why transfer learning and meta-learning generally cannot apply to the setting considered in CLLM . We provide empirical evidence on why the following dimensions are important (notably the point on the choice of downstream backbone). Specifically, we compare CLLM  with TabPFN (Hollmann et al., 2022), a few-shot learning method designed for small tabular problems (see Section C.9).
  1. 1.
    Access to external datasets: our problem setting in CLLM  assumes access to a single small training set $D_{\mathrm{train}}$, without access to any external/additional datasets. This mirrors the unique characteristics of low-to-middle income countries (LMICs), where data scarcity may be pervasive, hence making it unrealistic to assume that external datasets are available to the practitioner. In contrast, transfer learning and meta-learning make more stringent assumptions on the data requirements. (i) Transfer learning (Definition 3 in (Zhuang et al., 2020)): one typically assumes access to at least one additional source dataset $D_{\mathrm{source}}$, usually bigger than $D_{\mathrm{train}}$. This external $D_{\mathrm{source}}$ can then be used to pretrain a model, which is adapted using $D_{\mathrm{train}}$. (ii) Meta-learning (Hospedales et al., 2021): one assumes access to a set of $m$ tasks, which define a set of source datasets $\{D_{\mathrm{source}}^{(i)}\mid i\in[m]\}$. It is also worth noting that a common assumption in meta-learning is that the source datasets and $D_{\mathrm{train}}$ share the same feature space, restricting its applicability. To summarize, both these learning paradigms often rely on external data, an assumption not required in CLLM . We acknowledge that CLLM  can be seen as a form of transfer, in its broad definition, since it uses prior knowledge, with the LLM. However, a key point is that it does not require external/additional datasets.
  2. 2.
    Flexibility of the backbone model: since CLLM  is a data augmentation method for tabular data, it enables the practitioner to use any downstream model backbone, such as neural networks or tree-based methods (XGBoost, Random forest, CatBoost). This flexibility is important, as it allows the practitioner to choose the best-suited model for the task at hand. For example, one can use any tree-based method with CLLM , which is appealing given that tree-based methods are often preferred over neural networks for tabular data (Grinsztajn et al., 2022), (Shwartz-Ziv & Armon, 2022). In contrast, most approaches to transfer learning and meta-learning are not flexible as they traditionally require neural networks as backbone models. This stems from their methodology, where fine-tuning a pretrained model on $D_{\mathrm{train}}$ is a prevailing approach (Finn et al., 2017).
  3. 3.
    Ease of use: CLLM  is distinct in that it is easy to use, without the need for costly or complex operations, such as fine-tuning large pretrained models or using them at inference time, which may be impossible in LMICs, due to the cost associated to these operations.
@@
Active learning (Settles, 2009) and semi-supervised learning (van Engelen & Hoos, 2019; Chapelle et al., 2006) also operate under the premise of having access to plentiful unlabeled data and the capacity to interactively query labels. However, in our setting, considered learning with limited data does not inherently assume such capabilities, focusing instead on limited labeled data only. Furthermore, active learning primarily focuses on the iterative process of selecting data samples that, when labeled, are expected to most significantly improve the model’s performance. This selection is typically based on criteria such as uncertainty sampling which focuses on epistemic uncertainty (Mussmann & Liang, 2018; Houlsby et al., 2011; Kirsch et al., 2019; Nguyen et al., 2022). The primary objective is to minimize labeling effort while maximizing the model’s learning efficiency. Additionally, active learning would aims to label instances based on epistemic uncertainty where the model struggles to make accurate predictions, yet the samples themselves are correct. In contrast, CLLM leverage training dynamics based on aleatoric uncertainty and confidence and is designed to discard samples that might jeopardize the downstream accuracy. These samples can be considered to have inherent issues or are erroneous, such as being ”mislabeled”. To summarize, in active learning, epistemic uncertainty is used to identify data points that, if labeled, would yield the most significant insights for model training. In our approach, they serve to identify and exclude/filter data points that could potentially deteriorate the model’s performance.
@@
Self-supervised learning (Liu et al., 2021) leverages large amounts of unlabeled data to learn useful representations for downstream tasks. However, in our setting, considered learning with limited data does not inherently assume such access to vast amounts of unlabeled data.
@@
Data-centric AI. Ensuring high data quality is a critical but often overlooked problem in ML, where the focus is optimizing models (Sambasivan et al., 2021). Even when it is considered, the process of assessing datasets is adhoc or artisanal (Seedat et al., 2022b). However, the recent push of data-centric AI (Liang et al., 2022; Polyzotis & Zaharia, 2021; Zha et al., 2023; Seedat et al., 2023b) aims to develop systematic tools to curate existing datasets. Our work contributes to this nascent body of work (Seedat et al., 2023a) – presenting CLLM, which, to the best of our knowledge, is the first systematic data-centric framework looking at how we can tailor synthetic datasets (rather than real datasets) to downstream task use with data curation.
@@
Why Data Augmentation? Data augmentation is a flexible approach to address the low-data regime. An alternative might be to resort to a pretrained black-box model for classification, which could be for example via in-context learning for classification (Dong et al., 2022). However, such a solution is inadequate for several reasons, many of which would prevent real-world utility (e.g. in LMICs):
@@ B.1 Datasets
Adult Income (Asuncion & Newman, 2007) 30k 12 Finance
Compas (Angwin et al., 2016) 5k 13 Criminal justice
Covid-19 (Baqui et al., 2020) 7k 29 Healthcare/Medicine
CUTRACT Prostate (PCUK, 2019) 2k 12 Healthcare/Medicine
Drug (Fehrman et al., 2017) 2k 27 Healthcare/Medicine
MAGGIC (Pocock et al., 2013) 41k 29 Healthcare/Medicine
SEER Prostate (Duggan et al., 2016) 20k 12 Healthcare/Medicine
The private datasets are de-identified and used in accordance with the guidance of the respective data providers. We follow recommendations to use the Azure OpenAI service when using GPT-4 and GPT-3.5 models, where via the agreement we ensure the medical data is not sent for human review or stored, hence respecting the guidelines given by the dataset providers.
@@
    Open-source: Adult, Drug and Compas are widely used open-source datasets used in the tabular data literature. Adult and Drug are both UCI datasets that have been used in many papers, while Compas is part of OpenML (Vanschoren et al., 2013). Our reason for selecting them is that, despite them being open-source, they are highly reflective of domains in which we might be unable to collect many samples — hence in reality would often be in a low-data regime.
@@
GPT-4 and GPT-3.5 We access GPT-4 (OpenAI, 2023) and GPT-3.5-Turbo (Brown et al., 2020) through the API. We use a temperature of $0.9$.
GReaT. GReaT (Borisov et al., 2023) is a generative model which fine-tunes an LLM based on a training set. We use the implementation provided by authors.
Generative model based approaches. For the other baselines used in 3.1, we use the library SynthCity (Qian et al., 2023), using the defaults. We detail each next.
  • TVAE: this is a conditional Variational Auto Encoder (VAE) for tabular data and is based on Xu et al. (2019)
  • CTGAN: A conditional generative adversarial network which can handle tabular data and is based on Xu et al. (2019)
@@
  • TabDDPM: A diffusion model that can be universally applied to any tabular dataset, handles any type of feature and is based on Kotelnikov et al. (2022)
Traditional Data Augmentation. We use SMOTE (Chawla et al., 2002) which augments data by considering nearest neighbors and performing linear interpolations. We use the implementation provided by (Lemaître et al., 2017), and set the number of neighbors $k$ to $5$.